# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

### v2

@@
Abstract.
Scientific literature understanding is crucial for extracting targeted information and garnering insights, thereby significantly advancing scientific discovery. Despite the remarkable success of Large Language Models (LLMs), they face challenges in scientific literature understanding, primarily due to (1) a lack of scientific knowledge and (2) unfamiliarity with specialized scientific tasks.
@@ Abstract.
To develop an LLM specialized in scientific literature understanding, we propose a hybrid strategy that integrates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse scientific domain knowledge and enhance instruction-following capabilities for domain-specific tasks. In this process, we identify two key challenges: (1) constructing high-quality CPT corpora, and (2) generating diverse SFT instructions. We address these challenges through a meticulous pipeline, including PDF text extraction, parsing content error correction, quality filtering, and synthetic instruction creation. Applying this strategy, we present a suite of LLMs: SciLitLLM, specialized in scientific literature understanding. These models demonstrate promising performance on scientific literature understanding benchmarks. Specifically, the 7B model shows an average performance improvement of 3.6% on SciAssess and 10.1% on SciRIFF compared to leading LLMs with fewer than 15B parameters. Additionally, the 72B model, trained using QLoRA, achieves state-of-the-art performance among widely adopted open-source models.
Our contributions are threefold: (1) We present an effective framework that integrates CPT and SFT to adapt LLMs to scientific literature understanding, which can also be easily adapted to other domains. (2) We propose an LLM-based synthesis method to generate diverse and high-quality scientific instructions, resulting in a new instruction set – SciLitIns – for supervised fine-tuning in less-represented scientific domains. (3) SciLitLLM achieves promising performance improvements on scientific literature understanding benchmarks. We release the data processing codes [^1] https://github.com/dptech-corp/Uni-SMART and model weights [^2] https://huggingface.co/Uni-SMART/SciLitLLM.
Large Language Model, Pre-training, Supervised Fine-tuning, Scientific Literature Understanding
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing Methodologiesccs: Artificial Intelligenceccs: Natural Language Processingccs: Natural Language Generation
1. Introduction
@@
Figure 1. An example of scientific literature understanding in SciRIFF involves extracting accurate entities from a biomedicine paper. SciLitLLM-7B demonstrates sufficient scientific knowledge and instruction-following ability to accurately identify and extract these entities.
@@
Scientific literature understanding involves the systematic evaluation and interpretation of scientific texts and publications, to identify trends, extract targeted information, and garner insights (AI4Science and Quantum, 2023; Zheng et al., 2023), significantly contributing to scientific discovery. Concurrently, Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Yang et al., 2024; Touvron et al., 2023) have achieved remarkable success in natural language processing, prompting the development of domain-specific LLMs across various fields (Cui et al., 2024; Wu et al., 2023; Clusmann et al., 2023). However, recent studies (Wadden et al., 2024; Cai et al., 2024; Singh et al., 2023) indicate that LLMs face challenges when specializing in scientific literature understanding, particularly in context understanding and question answering. Take Figure 1 as an example, where the LLM is asked to understand the content of a biomedical research paper and then extract the targeted information. LLMs’ potential might be hindered by two major barriers: (1) a lack of scientific knowledge, which results in errors such as the missing important entities in Llama-3-7B (AI@Meta, 2024), and (2) unfamiliarity with scientific tasks, leading to the inability of Galactica-6.7B (Taylor et al., 2022) to follow task instructions accurately.
@@
To make LLMs specialized in science-relevant tasks, existing studies mostly adopt two strategies, as illustrated in Figure 2: (1) Fine-tuning with scientific instructions (Wadden et al., 2024; Zhang et al., 2024b; Singhal et al., 2022; Huang et al., 2023). A general-purpose LLM is fine-tuned with collected domain-specific instructions to adapt it to science-relevant tasks. However, instruction fine-tuning alone is insufficient to imbue the models with comprehensive scientific knowledge. (2) Pre-training on scientific corpora (Beltagy et al., 2019; Zeng et al., 2022; Taylor et al., 2022). This approach involves training models on vast scientific corpora. While this method equips LLMs with domain knowledge, the lack of instruction-tuning confines them to solving relevant tasks. Moreover, it is hampered by substantial computational costs and data requirements (Ling et al., 2023; Yang et al., 2023). To address these obstacles while balancing efficiency, we propose a hybrid strategy that incorporates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse domain knowledge and enhance domain-specific instruction-following capabilities.
Figure 2. Comparison of strategies to adapt LLMs to scientific tasks. Previous approaches typically either fine-tune a general LLM with scientific instructions or pre-train an LLM on extensive scientific corpora. We propose a combined method of both CPT and SFT.
However, as illustrated in Figure 3, developing a scientific literature understanding model using this CPT and SFT pipeline presents two critical requirements:
@@
  • High-quality CPT Corpora. Scientific corpora, predominantly in PDF format such as textbooks and research papers, are not directly digestible for LLM training. Converting these documents to text using tools like PyPDF2333https://pypdf2.readthedocs.io often introduces formatting and syntax errors, degrading corpus quality. Worse still, scientific documents often contain segments that contribute little information (e.g., references), necessitating quality control to filter them out. See the first row in Figure 3 for a comparison of high- and low-quality CPT texts.
  • Diverse Scientific Instructions. Effective instruction following for scientific literature understanding requires a large, high-quality, and diverse set of task-related instructions. However, to the best of our knowledge, there is a scarcity of well-designed instruction datasets for scientific literature understanding, and hiring human annotators to curate such a dataset from scratch is prohibitively expensive (Erdmann et al., 2019; Qiu et al., 2023). See the second row in Figure 3 for an illustration of high- and low-quality instructions.
@@
  • In the CPT stage for domain knowledge injection, we start with an extensive in-house corpus consisting of 73k textbooks and 625k academic papers in the scientific field, all in PDF format. Initially, we leverage PyPDF2, a widely used open-source PDF parsing tool, to extract raw texts from these documents. We then employ a moderate yet powerful model, Llama3-7B-Instruct (AI@Meta, 2024), to correct the format and spelling errors introduced by PDF parsing (cf. Section 3.1.1). Subsequently, we train a small text quality classifier to score the corpus and filter out texts of low educational value [^4] Phi models (Gunasekar et al., 2023; Li et al., 2023a; Abdin et al., 2024) propose to determine the quality of a pre-training text by its educational value for a student whose goal is to learn basic domain concepts. in the scientific field (cf. Section 3.1.2). These two simple yet effective textual refinement and quality control measures ensure the high quality of our CPT corpus, culminating 12.7 billion tokens for CPT via the Qwen2 (Yang et al., 2024) tokenizer.
@@
Figure 3. Examples of high and low-quality CPT text and SFT instructions. Scores are labeled by the CPT and SFT quality filter (cf. Section 3.1.2 &  3.2.2), ranging from 0 to 5. Higher scores indicate better quality.
@@
Having established such high-quality datasets, we apply the CPT-SFT integration strategy on a general-purpose LLM – Qwen2 (Yang et al., 2024) and obtain SciLitLLM in two scales: a 7B (full parameter training) and a 4-bit quantized 72B (QLoRA (Dettmers et al., 2023) parameter-efficient training) model. Evaluations on benchmarks of scientific literature understanding demonstrate the effectiveness of our strategy. We observe promising performance enhancements, with an average improvement of 3.6% on SciAssess (Cai et al., 2024) and 10.1% on SciRIFF (Wadden et al., 2024), compared to the leading LLMs under 15B parameters. Notably, SciLitLLM-7B even outperforms Llama3 and Qwen2 with 70B parameters on SciRIFF. Additionally, SciLitLLM-72B, with QLoRA training, achieves leading results on both benchmarks, surpassing other open-source LLMs. Further ablation studies demonstrate the effectiveness of each module in our pipeline.
In summary, our contributions are threefold:
@@
  • We devise an effective and comprehensive pipeline to adapt general LLMs to a specific domain – scientific literature understanding. It combines continual pre-training (CPT) and supervised fine-tuning (SFT), to enhance scientific knowledge base and instruction-following capabilities for specialized domain tasks.
  • We propose a novel domain instruction synthesis method to curate instructions for scientific literature understanding, resulting in a new dataset – SciLitIns .
  • SciLitLLM, trained through the proposed pipeline, outperforms leading open-source LLMs on scientific literature understanding. It has been deployed in real-world application scenarios and has demonstrated promising performance. Code and data will be released after administrative procedure.
2. Related Works
In this section, we provide a review of literature related to continual pre-training, supervised fine-tuning in the scientific domain, and LLMs for scientific literature understanding.
2.1. Knowledge Injection via Continual Pre-training
Pre-training a language model is usually conducted on a large corpus of textual data to learn the statistical properties of language (Radford et al., 2019; Brown et al., 2020). Formally, given a sequence of textual tokens $x=(x_{1},x_{2},...,x_{T})$, an LLM parameterized by $\theta$ performs autoregressive language modeling (ALM) by predicting the next token in the sequence given the previous tokens:
(1) $\mathcal{L}_{\text{ALM }}(\theta)=\sum_{t=1}^{T}\log P_{\theta}\left(x_{t}\mid x_{<t}\right).$
@@ 2.1. Knowledge Injection via Continual Pre-training
To further inject domain knowledge into a general LLM after pre-training, researchers engage in continual pre-training (CPT) on high-quality domain-specific corpora (Jin et al., 2022; Sun et al., 2020), sometimes combined with general corpora. This process enhances the model’s fundamental understanding abilities in specific downstream domains while mitigating catastrophic forgetting of general knowledge (Wu et al., 2022; Ke et al., 2023; Mehta et al., 2023). See the comprehensive study (Gupta et al., 2023) for different warm-up strategies for CPT. Additionally, the CPT corpora can be augmented by transforming them into an instruction-response format (Cheng et al., 2023, 2024). Furthermore, the scaling law (Hoffmann et al., 2022) of domain-specific CPT (Que et al., 2024) is explored to determine the optimal mix of data. However, these studies primarily focus on training dynamics and data recipes, leaving the pre-processing for scientific data, especially raw PDF files, largely unexplored. Exhibiting such steps is essential for generating high-quality domain corpora and effectively injecting domain knowledge, representing a significant challenge for practitioners.
2.2. Domain Adaptation via Supervised Fine-tuning
Supervised fine-tuning (SFT) modifies a pre-trained language model to follow specific instructions or perform designated tasks by fine-tuning it on a targeted, task-specific dataset (Raffel et al., 2020; Wei et al., 2021). Let $\mathcal{D}_{\text{fine}}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ represent the fine-tuning dataset, where $x_{i}$ is an input sequence (instruction) and $y_{i}$ is the corresponding output sequence. The fine-tuning objective can be expressed as:
(2) $\mathcal{L}_{\text{fine-tune }}(\phi)=\sum_{i=1}^{M}\log P_{\phi}\left(y_{i}\mid x_{i}\right).$
@@ 2.2. Domain Adaptation via Supervised Fine-tuning
During fine-tuning, the model adjusts its parameters $\phi$ to better fit the task-specific data, typically involving parameter-efficient (Hu et al., 2021; Dettmers et al., 2023) or full parameter training (Raffel et al., 2020; Wei et al., 2022a). Applying SFT to a general LLM for specific domain adaptation has demonstrated effectiveness in various fields: in medicine (Clusmann et al., 2023), corpora of medical literature and clinical notes are used; in law (Cui et al., 2024), legal documents and case law are compiled; and in finance (Wu et al., 2023), financial reports and market data are utilized. In the scientific domain, several studies have specialized LLMs for scientific tasks, often necessitating the construction of a substantial domain-specific dataset with SFT. For example, SciGLM (Zhang et al., 2024a) leverages existing LLMs to generate step-by-step reasoning for unlabelled scientific instructions. ChemLLM (Zhang et al., 2024b), a more specified LLM in the chemistry field, collects structured chemical data from a vast selection of online databases and transforms this structured data into a question-answering format for SFT. SciRIFF (Wadden et al., 2024) converts existing literature understanding datasets into natural language input-output pairs suitable for instruction-tuning using pre-defined templates. However, benchmark studies (Feng et al., 2024; Cai et al., 2024) indicate that SFT alone may not provide adequate scientific knowledge to excel in relevant tasks. This suggests the need for a more comprehensive approach that combines domain knowledge infusion with instruction-following enhancements.
Figure 4. The pipeline of SciLitLLM consists of two key stages: continual pre-training (CPT) for scientific knowledge injection and supervised fine-tuning (SFT) for scientific instruction following. Specifically, the CPT stage involves several modules: PDF parsing, format & grammar correction (cf. Section 3.1.1), and quality filtering (cf. Section  3.1.2) modules. These modules ensure the model is equipped with high-quality scientific domain knowledge. The SFT stage includes LLM-based instruction generation (cf. Section 3.2.1) and instruction quality control (cf. Section 3.2.2) measures. These steps are designed to fine-tune the model’s ability to follow scientific instructions accurately and effectively.
2.3. LLMs for Scientific Literature Understanding
@@
3. Method
@@
In this section, we discuss the details of our proposed pipeline (cf. Figure 4): continual pre-training for scientific knowledge injection (cf. Section 3.1) and supervised fine-tuning for scientific tasks enhancement (cf. Section 3.2).
3.1. CPT for Scientific Knowledge Injection
@@
What are high-quality pre-training corpora? Researchers (Gunasekar et al., 2023; Li et al., 2023a; Abdin et al., 2024) suggest that language models benefit from corpora that possess the same qualities as an exemplary textbook for human learners: clarity, self-containment, instructiveness, and balance. These characteristics ensure that the material is not only comprehensible but also informative and comprehensive, providing a solid foundation for knowledge acquisition. Over the past decades, the efforts of scientists and educators have resulted in a wealth of high-quality scientific textbooks and research papers, which serve as invaluable resources for learning and teaching. Recognizing this, we have curated a substantial collection of over 73,000 textbooks and 625,000 research papers within the scientific domain, ensuring all documents are copyright-compliant. To inject their rich scientific knowledge into a general LLM, we perform continual pre-training (CPT) on these high-quality textbooks and papers. This process equips the model with a robust scientific knowledge base, thereby paving the way for developing a specialized LLM tailored for scientific literature understanding.
@@
However, we still face two practical obstacles when dealing with those textbooks and research papers: (1) Formatting and syntax errors. Most textbooks and research paper documents are in PDF format, which is not directly digestible by LLMs. Consequently, we need to transform them into plain text. Converting these documents using tools like PyPDF2 often introduces formatting and syntax errors, which degrade the quality of the corpus. (2) Corpus quality control. Despite their overall high quality, textbooks and research papers also contain segments with little useful information, such as references and garbled text introduced during the PDF parsing process. Given the large scale of the pre-training corpora, an effective and computation-efficient quality control measure is essential.
To tackle these obstacles, we devised the following modules of format & grammar correction and CPT quality filter:
3.1.1. Format & Grammar Correction

### v3

@@
Abstract
@@
Scientific literature understanding is crucial for extracting targeted information and garnering insights, thereby significantly advancing scientific discovery. Despite the remarkable success of Large Language Models (LLMs), they face challenges in scientific literature understanding, primarily due to (1) a lack of scientific knowledge and (2) unfamiliarity with specialized scientific tasks. To develop an LLM specialized in scientific literature understanding, we propose a hybrid strategy that integrates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse scientific domain knowledge and enhance instruction-following capabilities for domain-specific tasks. In this process, we identify two key challenges: (1) constructing high-quality CPT corpora, and (2) generating diverse SFT instructions. We address these challenges through a meticulous pipeline, including PDF text extraction, parsing content error correction, quality filtering, and synthetic instruction creation. Applying this strategy, we present a suite of LLMs: SciLitLLM, specialized in scientific literature understanding. These models demonstrate promising performance on scientific literature understanding benchmarks. Our contributions are threefold: (1) We present an effective framework that integrates CPT and SFT to adapt LLMs to scientific literature understanding, which can also be easily adapted to other domains. (2) We propose an LLM-based synthesis method to generate diverse and high-quality scientific instructions, resulting in a new instruction set – SciLitIns – for less-represented scientific domains. (3) SciLitLLM achieves promising performance in scientific literature understanding benchmarks. We release the data processing codes [^1] https://github.com/dptech-corp/Uni-SMART/tree/main/SciLitLLM, model weights [^2] https://huggingface.co/Uni-SMART/SciLitLLM-1.5 and the synthetic dataset [^3] https://huggingface.co/datasets/Uni-SMART/SciLitIns.
1 Introduction
Figure 1: Average scores on SciRIFF of models with varying parameter sizes.
@@ 1 Introduction
Figure 2: An example of scientific literature understanding in SciRIFF. It involves extracting accurate entities from a biomedicine paper. SciLitLLM-7B demonstrates sufficient scientific knowledge and instruction-following ability to accurately identify and extract these entities.
@@
Scientific literature understanding involves the systematic evaluation and interpretation of scientific texts and publications, to identify trends, extract targeted information, and garner insights (AI4Science & Quantum, 2023; Zheng et al., 2023), significantly contributing to scientific discovery. Concurrently, Large Language Models (LLMs) (OpenAI, 2023; Team, 2024; Touvron et al., 2023) have achieved remarkable success in natural language processing, prompting the development of domain-specific LLMs across various fields (Cui et al., 2024; Wu et al., 2023; Clusmann et al., 2023). However, recent studies (Wadden et al., 2024; Cai et al., 2024; Singh et al., 2023) indicate that LLMs face challenges when specializing in scientific literature understanding, particularly in context understanding and question answering. Take Figure 2 as an example, where the LLM is asked to understand the content of a biomedical research paper and then extract the targeted information. LLMs’ potential might be hindered by two major barriers: (1) a lack of scientific knowledge, which results in errors such as the missing important entities in Llama-3.1-8B (AI@Meta, 2024), and (2) unfamiliarity with scientific tasks, leading to the inability of Galactica-6.7B (Taylor et al., 2022) to follow task instructions accurately.
@@
To make LLMs specialized in science-relevant tasks, existing studies mostly adopt two strategies, as illustrated in Figure 4: (1) Fine-tuning with scientific instructions (Wadden et al., 2024; Zhang et al., 2024b; Singhal et al., 2022). A general-purpose LLM is fine-tuned with collected domain-specific instructions to adapt it to science-relevant tasks. However, instruction fine-tuning alone is insufficient to imbue the models with comprehensive scientific knowledge. (2) Pre-training on scientific corpora (Beltagy et al., 2019; Zeng et al., 2022; Taylor et al., 2022). This approach involves training models on vast scientific corpora. While this method equips LLMs with domain knowledge, the lack of instruction-tuning confines them to solving relevant tasks. Moreover, it is hampered by substantial computational costs and data requirements (Ling et al., 2023; Yang et al., 2023). To address these obstacles while balancing efficiency, we propose a hybrid strategy that incorporates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse domain knowledge and enhance domain-specific instruction-following capabilities.
@@
However, as illustrated in Figure 4, developing a scientific literature understanding model using this CPT and SFT pipeline presents two critical requirements: (1) High-quality CPT Corpora. Scientific corpora, predominantly in PDF format such as textbooks and research papers, are not directly digestible for LLM training. Converting these documents to text with PDF parsing tools introduces formatting and syntax errors, degrading corpus quality. Worse still, scientific documents often contain segments that contribute little information (e.g., references), necessitating quality control to filter them out. See the first row in Figure 4 for a comparison of high- and low-quality CPT texts. (2) Diverse Scientific Instructions. Effective instruction following for scientific literature understanding requires a large, high-quality, and diverse set of task-related instructions. However, to the best of our knowledge, there is a scarcity of well-designed instruction datasets for scientific literature understanding, and hiring human annotators to curate such a dataset from scratch is prohibitively expensive (Erdmann et al., 2019; Qiu et al., 2023). See the second row in Figure 4 for an illustration of high- and low-quality instructions.
@@
In the CPT stage for domain knowledge injection, we start with an extensive in-house corpus consisting of 73k textbooks and 625k academic papers in the scientific field, all in PDF format. Initially, we leverage PyPDF2444https://pypdf2.readthedocs.io, a widely used open-source PDF parsing tool, to extract raw texts from these documents. We then employ a moderate yet powerful model, Llama3-7B-Instruct, to correct the format and spelling errors introduced by PDF parsing (cf. Section 3.1.1). Subsequently, we train a small text quality classifier to score the corpus and filter out texts of low educational value [^5] Phi models (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024) propose to determine the quality of a pre-training text by its educational value for a student whose goal is to learn basic domain concepts. in the scientific field (cf. Section 3.1.2). These two simple yet effective textual refinement and quality control measures ensure the high quality of our CPT corpus, culminating 12.7 billion tokens for CPT via the Qwen2.5 tokenizer.
@@
Figure 3: Comparison of strategies to adapt LLMs to scientific tasks.
Figure 4: Examples of high and low-quality CPT text and SFT instructions.
@@
Having established such high-quality datasets, we apply the CPT-SFT integration strategy on a general-purpose LLM – Qwen2.5 (Team, 2024) and obtain SciLitLLM of two scales: 7B and 14B. Evaluations on benchmarks of scientific literature understanding demonstrate the effectiveness of our strategy. We observe promising performance enhancements, with an average improvement of 4.0% on SciAssess (Cai et al., 2024) and 10.1% on SciRIFF (Wadden et al., 2024), compared to the leading LLMs under 10B parameters. Notably, SciLitLLM-7B even outperforms Llama3.1 and Qwen2.5 with 70B parameters on SciRIFF. Additionally, SciLitLLM-14B achieves leading results on both benchmarks, surpassing other open-source LLMs. Further ablation studies demonstrate the effectiveness of each module in our pipeline.
@@
In summary, our contributions are threefold: (1) We devise an effective and comprehensive pipeline to adapt general LLMs to a specific domain – scientific literature understanding. It combines continual pre-training (CPT) and supervised fine-tuning (SFT), to enhance scientific knowledge base and instruction-following capabilities for specialized domain tasks. (2) We propose a novel domain instruction synthesis method to curate instructions for scientific literature understanding, resulting in a new dataset – SciLitIns. (3) SciLitLLM, trained through the proposed pipeline, outperforms leading open-source LLMs on scientific literature understanding.
2 Related Works
2.1 Knowledge Injection via Continual Pre-training
@@
Pre-training a language model is usually conducted on a large corpus of textual data to learn the statistical properties of language (Radford et al., 2019; Brown et al., 2020). To further inject domain knowledge into a general LLM after pre-training, researchers engage in continual pre-training (CPT) on high-quality domain-specific corpora (Jin et al., 2022; Sun et al., 2020), sometimes combined with general corpora. This process enhances the model’s fundamental understanding abilities in specific downstream domains while mitigating catastrophic forgetting of general knowledge (Wu et al., 2022; Ke et al., 2023; Mehta et al., 2023). See the comprehensive study (Gupta et al., 2023) for different warm-up strategies for CPT. Additionally, the CPT corpora can be augmented by transforming them into an instruction-response format (Cheng et al., 2023; 2024). Furthermore, the scaling law (Hoffmann et al., 2022) of domain-specific CPT (Que et al., 2024) is explored to determine the optimal mix of data. However, these studies primarily focus on training dynamics and data recipes, leaving the pre-processing for scientific data, especially raw PDF files, largely unexplored. Exhibiting such steps is essential for generating high-quality domain corpora and effectively injecting domain knowledge, representing a significant challenge for practitioners.
2.2 Domain Adaptation via Supervised Fine-tuning
@@
Supervised fine-tuning (SFT) modifies a pre-trained language model to follow specific instructions or perform designated tasks by fine-tuning it on a targeted, task-specific dataset (Raffel et al., 2020; Wei et al., 2021). Applying SFT to a general LLM for specific domain adaptation has demonstrated effectiveness in various fields: in medicine (Clusmann et al., 2023), corpora of medical literature and clinical notes are used; in law (Cui et al., 2024), legal documents and case law are compiled; and in finance (Wu et al., 2023), financial reports and market data are utilized. In the scientific domain, several studies have specialized LLMs for scientific tasks, often necessitating the construction of a substantial domain-specific dataset with SFT. For example, SciGLM (Zhang et al., 2024a) leverages existing LLMs to generate step-by-step reasoning for unlabelled scientific instructions. ChemLLM (Zhang et al., 2024b), a more specified LLM in the chemistry field, collects structured chemical data from a vast selection of online databases and transforms this structured data into a question-answering format for SFT. SciRIFF (Wadden et al., 2024) converts existing literature understanding datasets into natural language input-output pairs suitable for instruction-tuning using pre-defined templates. However, benchmark studies (Feng et al., 2024; Cai et al., 2024) indicate that SFT alone may not provide adequate scientific knowledge to excel in relevant tasks. This suggests the need for a more comprehensive approach that combines domain knowledge infusion with instruction-following enhancements.
2.3 LLMs for Scientific Literature Understanding
@@
3 Method
Figure 5: The pipeline of SciLitLLM consists of two key stages: continual pre-training (CPT) for scientific knowledge injection and supervised fine-tuning (SFT) for scientific instruction following. Specifically, the CPT stage involves several modules: PDF parsing, format & grammar correction (cf. Section 3.1.1), and quality filtering (cf. Section  3.1.2) modules. These modules ensure the model is equipped with high-quality scientific domain knowledge. The SFT stage includes LLM-based instruction generation (cf. Section 3.2.1) and instruction quality control (cf. Section 3.2.2) measures. These steps are designed to fine-tune the model’s ability to follow scientific instructions accurately and effectively.
@@ 3 Method
In this section, we present the details of our proposed pipeline (cf. Figure 5): continual pre-training for scientific knowledge injection and supervised fine-tuning for scientific tasks enhancement.
3.1 CPT for Scientific Knowledge Injection
@@
What are high-quality pre-training corpora? Researchers (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024) suggest that language models benefit from corpora that possess the same qualities as an exemplary textbook for human learners: clarity, self-containment, instructiveness, and balance. Recognizing the wealth of high-quality scientific textbooks and research papers over the past decades, we have curated a substantial collection of over 73,000 textbooks and 625,000 research papers within the scientific domain, ensuring all documents are copyright-compliant.
@@
However, we still face two practical obstacles when dealing with those textbooks and research papers: (1) Formatting and syntax errors. Most textbooks and research paper documents are in PDF format, which is not directly digestible by LLMs. Converting these documents using tools like PyPDF2 often introduces formatting and syntax errors, which degrade the quality of the corpus. (2) Corpus quality control. Despite their overall high quality, textbooks and research papers also contain segments with little useful information, such as references and garbled text introduced during the PDF parsing process. To tackle these obstacles, we devised the following modules: format & grammar correction and CPT quality filter.
3.1.1 Format & Grammar Correction